Purchases Budget

Rest Inn provides four-star accommodations for the vacation traveler. It is located just off a major interstate freeway. There are three other competing motels at the same exit. Hotel management is preparing its budget for the busiest three-month period of the year, June through August. To differentiate themselves from the competition, Rest Inn provides a complimentary spa package for each guest stay that includes specialty shampoo, conditioner, soap, lotion, toothpaste, lavender essential oil, slippers, earplugs, and a sleep mask. Rest Inn purchases the spa package from a local vendor that puts the Rest Inn private label on the products. The hotel follows a policy of purchasing enough spa packages to ensure that 40% of next month’s bookings are in the current month’s ending inventory. Rest Inn’s sales budget for guest stays for the next quarter is as follows:

June 2,200 guest stays
July 2,400 guest stays
August 2,000 guest stays

How many spa packages should Rest Inn budget for purchase in July?

Answer spa packages
Answer #1
Guest stays for July 2400
Add: Desired ending inventory 800
Total needed 3200
Less: Beginning inventory -960
Purchases in July 2240

Desired ending inventory= August guest stays*40%= 2000*40%= 800

Beginning inventory= July guest stays*40%= 2400*40%= 960

How many spa packages should Rest Inn budget for purchase in July?

2240 packages